# IndiSeek learns information-guided disentangled representations

### v1

@@ A.4.1 Experimental results with varying $\lambda$
A.4.2 Multi-task performance on MIMIC dataset
@@
Table 8: Multi-task performance on MIMIC (averaged over 10 seeds).
@@
0 77.79% 77.48% 77.21% 77.21%
1 91.32% 91.28% 91.26% 91.13%
2 70.83% 70.64% 70.76% 70.40%
3 66.64% 66.58% 66.76% 66.49%
4 70.98% 70.94% 70.71% 70.62%
5 72.16% 71.94% 72.07% 71.83%
6 85.87% 85.91% 85.89% 85.73%
7 65.83% 65.87% 65.84% 64.78%
8 65.84% 65.83% 65.42% 65.29%
9 73.15% 73.17% 72.77% 72.50%
10 99.60% 99.60% 99.60% 99.60%
11 89.14% 89.14% 89.14% 89.14%
12 80.93% 80.93% 80.93% 80.93%
13 96.69% 96.68% 96.69% 96.68%
14 69.34% 69.48% 69.19% 69.03%
15 90.98% 90.98% 90.98% 90.98%
16 96.84% 96.84% 96.84% 96.84%
17 62.26% 61.88% 61.90% 61.33%
18 60.92% 60.53% 60.57% 60.80%
19 69.56% 69.43% 69.50% 69.33%
Average 77.83% 77.76% 77.70% 77.53%
In Figure 8, IndiSeek leads the performance in most tasks. Moreover, in many real-world scenarios, downstream tasks are unknown in advance, or learned representations are expected to handle hundreds of tasks simultaneously. In this case, the task-relevant optimality of augmentations could be barely met in Liang et al., (2023), which highlights the flexibility and practicality of task-agnostic disentangled learning.

### v3

@@ A.4.1 Experimental results with varying $\lambda$
A.4.2 Training time comparison
We compare the computational efficiency of different methods by measuring the training time required to compute modality-specific features. Specifically, we record the time for training the disentangled representations for both modalities (text and vision) after the initial CLIP feature extraction. All methods are trained with consistent configurations: we use a batch size of 128, a maximum of 2000 epochs with early stopping, Adam optimizer with learning rate $10^{-4}$ and weight decay $10^{-4}$, and Transformer-based encoders to map raw sequences to a $50$-dimensional latent space. The training is performed on a single GPU, and we measure the wall-clock time for the complete training of both modality-specific encoders.
Table 9 presents the training time comparison across datasets, averaged over 10 random seeds. The results show that all three methods have comparable computational costs, with differences primarily stemming from dataset sizes and sequence lengths. IndiSeek demonstrates competitive efficiency while achieving superior accuracy (as shown in Table 1), making it a practical choice for multimodal learning tasks.
A.4.3 Ablation studies and comparison with CoMM
To understand the contribution of individual components in IndiSeek, we conduct an ablation study on the MOSI dataset. Specifically, we compare IndiSeekmethod against IndiSeek0, a variant that removes the modality-switching mechanism and uses a fixed pairing between raw inputs and CLIP embeddings. We also include a comparison with CoMM (dufumier2024align). Table 10 presents the results of this analysis. CoMM achieves competitive performance at 69.56%, showing that collaborative learning approaches can be effective for multimodal representation learning.
Table 10: MOSI ablation study results with varying $\lambda$ (averaged over 10 seeds). All values are percentages.
$\lambda$ 0.01 0.1 1.0 10.0 100.0 1000.0 max over $\lambda$
IndiSeek 67.11(0.85) 69.61(0.55) 70.09(0.80) 66.63(0.56) 66.43(0.50) 66.06(0.72) 70.09(0.80)
IndiSeek0 66.24(0.43) 69.21(0.48) 69.59(0.27) 66.57(0.49) 65.73(0.35) 65.07(0.21) 69.59(0.27)
CoMM 68.34(0.79) 68.69(0.65) 69.56(0.47) 67.14(0.37) 66.79(0.59) 66.60(0.44) 69.56(0.47)
FactorizedCL 66.76(0.41) 65.07(0.65) 66.12(0.47) 64.26(0.76) 65.07(0.41) 65.12(0.63) 67.11(0.24)
InfoDisen 65.57(0.59) 65.31(0.89) 65.83(0.71) 65.22(0.76) 64.84(0.89) 67.52(0.44) 67.52(0.44)
CLIP 67.14(0.87) 67.14(0.87) 67.14(0.87) 67.14(0.87) 67.14(0.87) 67.14(0.87) 67.14(0.87)
A.4.4 Multi-task performance on MIMIC dataset
@@
Table 11: Multi-task performance on MIMIC dataset (max over $\lambda$, averaged over 10 seeds). All values are percentages.
@@
0 77.73(0.06) 77.76(0.10) 77.34(0.04) 77.07(0.18)
1 91.31(0.03) 91.30(0.02) 91.30(0.02) 91.12(0.07)
2 70.76(0.19) 70.67(0.11) 70.62(0.24) 70.22(0.09)
3 67.04(0.10) 67.01(0.08) 66.78(0.12) 66.89(0.11)
4 71.04(0.22) 70.86(0.32) 70.87(0.20) 70.81(0.07)
5 72.22(0.08) 72.03(0.05) 72.11(0.09) 71.97(0.09)
6 86.00(0.06) 86.17(0.10) 85.94(0.11) 85.98(0.07)
7 66.04(0.18) 65.81(0.35) 65.58(0.34) 64.50(0.42)
8 65.66(0.07) 65.84(0.05) 65.50(0.19) 65.04(0.12)
9 73.09(0.17) 73.16(0.09) 72.75(0.19) 72.44(0.19)
10 99.60(0.00) 99.60(0.01) 99.60(0.00) 99.60(0.00)
11 89.14(0.00) 89.14(0.00) 89.14(0.00) 89.14(0.00)
12 80.93(0.00) 80.94(0.02) 80.93(0.01) 80.93(0.00)
13 96.69(0.00) 96.69(0.02) 96.69(0.01) 96.69(0.00)
14 69.43(0.05) 69.21(0.17) 69.13(0.13) 68.99(0.02)
15 90.98(0.00) 90.98(0.00) 90.98(0.00) 90.98(0.00)
16 96.84(0.00) 96.84(0.00) 96.84(0.00) 96.84(0.00)
17 62.58(0.17) 61.83(0.20) 61.72(0.14) 61.50(0.14)
18 60.99(0.19) 60.36(0.14) 60.40(0.34) 60.71(0.09)
19 69.71(0.11) 69.30(0.12) 69.46(0.14) 69.25(0.16)
Average 77.89 77.82 77.74 77.55
In Figure 11, IndiSeek leads the performance in most tasks. Moreover, in many real-world scenarios, downstream tasks are unknown in advance, or learned representations are expected to handle hundreds of tasks simultaneously. In this case, the task-relevant optimality of augmentations could be barely met in liang2023factorized, which highlights the flexibility and practicality of task-agnostic disentangled learning.